Find three consecutive odd integers such that three times the second minus the third is 11 more than the first.
I set the equation up as follows:
3(x + 2) - x + 4 = x + 11
Thanks for your help

Answer by checkley71(8403)

3(X+2)-(X+4)=X+11
3X+6-X-4=X+11
3X-X-X=11-6+4
X=9 ANSWER FOR THE SMALLEST NUMBER
THE OTHER NUMBERS ARE 11 & 13.
PROOF
3*11-13=9+11
33-13=20
20=20